Circular No. 6269

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 G. Pooley, Mullard Radio Astronomy Observatory, Cambridge, writes:

“Observations of GRS 1915+105 with the Ryle Telescope at 15 GHz show evidence for variations in flux density with a period close to 38 min. During these oscillations, the mean flux is in the range 10–30 mJy and the amplitude varies typically by a factor of 2. Oscillations were observed on Sept. 8, Nov. 15, 16, 17, 29, 30, and Dec. 1, but none was apparent on any of 16 observation days during October. It is not yet possible to say whether the oscillations are coherent in phase from one day to the next. It is important to search for variations at other wavelengths.”

SATELLITES OF SATURN B. Sicardy, Paris Observatory; J.-L. Beuzit, European Southern Obser-

vatory (ESO), Garching; F. Poulet, Paris Observatory; and P. Prado, ESO, La Silla, report: “During a 6-night period spanning the Aug. 10 ring-plane crossing, we obtained 60-s exposures of Saturn’s rings using the ESO 3.6m telescope (+ Adonis adaptive optics camera + short K filter; bandpass centered at 2.16 μm). Besides Saturn X (Janus), Saturn XI (Epimetheus), and Saturn XVII (Pandora), observed at their expected locations, an unresolved object is detected on two occasions, at maximum eastern and western elongations, respectively. It then coincides with the location of the F ring. No conspicuous point-like object is visible there on good-quality images taken at other times. The object is visible on Aug. 9.4388 and 9.4493 UT, located 21′′.6 and 21′′.7 eastward of Saturn’s center, respectively. During Aug. 10.3462–10.3710, it appears 21′′.9 westward of the planet’s center on seven images. Typical astrometric accuracy is estimated as 0′′.2; within this accuracy, these positions agree with the ephemerides of the object S/1995 S6, discovered with the Hubble Space Telescope (HST) a few hours after our last observations (cf. IAUC 6243). No other ESO images are available at the times when S/1995 S6 should have been visible. The ESO detections may be prediscovery images of S/1995 S6. However, comparison of the ESO object brightness with Saturn III (Tethys) yields a consistent K magnitude of 14.7 ± 0.2. This indicates a geometric albedo intrinsically five times higher in our short K filter (2.16 μm) than in the HST methane band (0.89 μm), from which the value V = 17.8 is derived (ibid.). This large color effect is not yet explained.”

1995 December 4 Daniel W. E. Green